OpenStudy (anonymous):

Which of the following best defines 3 to the power of 2 over 3

OpenStudy (jim766):

3 * 3 divide by 3

OpenStudy (xguardians):

\[\sqrt[3]{3^{2}}\]

OpenStudy (anonymous):

theres more to it ...the answers are Square root of 9 Cube root of 9 Cube root of 3 Square root of 3

OpenStudy (jim766):

first, what is (3*3) / 3

OpenStudy (anonymous):

6/3 ??

OpenStudy (jim766):

wait is 3 to the power of a fraction 2/3

OpenStudy (anonymous):

2/3=1/3 right ?

OpenStudy (anonymous):

it would be 2

OpenStudy (jim766):

if the question is what is 3 ^ 2/3 it is the cube root of 9

OpenStudy (jim766):

3^2/3 is the same as cube root of 3^2

OpenStudy (jim766):

the 2nd choice

OpenStudy (anonymous):

oh. im taking algebra online and im bad at it so i dont know that thank you so much. can you help me on few more pleasse?